# Version of Tuesday $24^{\text{th}}$ June, 2025

### v1

@@ 3.2. Sufficient conditions: Conjugate points
As it is well understood, satisfying the Euler-Lagrange equation (9) gives an extremal path, but it is only a necessary condition for a minimizer. The present concern is to obtain sufficient, and verifiable, conditions for having obtained a minimizer in (6). The original work of Gelfand, [10], lays the groundwork for providing sufficient conditions, though both [10] as well as later books on the topic (e.g., see [11, p.19] and [13, p.54]) derive sufficient conditions under the assumption that the matrix ${\mathcal{L}}_{{\mathbf{x}}\dot{\mathbf{x}}}$ is symmetric, which is not satisfied in our case. For this reason, here we provide the full derivation of sufficient conditions, by looking at the 2nd variation of the functional $\mathcal{L}$} along a trajectory ${\mathbf{x}}(t)$ that solves the Euler-Lagrange equation.
@@
(15) $\begin{split}\delta^{2}{\mathcal{L}}=\frac{1}{2}&\int_{0}^{1}\left({\mathbf{y}}^{T}{\mathcal{L}}_{{\mathbf{x}}{\mathbf{x}}}{\mathbf{y}}+\dot{\mathbf{y}}^{T}{\mathcal{L}}_{\dot{\mathbf{x}}\dot{\mathbf{x}}}\dot{\mathbf{y}}+\dot{\mathbf{y}}^{T}{\mathcal{L}}_{\dot{\mathbf{x}}{\mathbf{x}}}{\mathbf{y}}+{\mathbf{y}}^{T}{\mathcal{L}}_{{\mathbf{x}}\dot{\mathbf{x}}}\dot{\mathbf{y}}\right)dt\\ :=\frac{1}{2}&\int_{0}^{1}\left({\mathbf{y}}^{T}A{\mathbf{y}}+\dot{\mathbf{y}}^{T}C\dot{\mathbf{y}}+\dot{\mathbf{y}}^{T}B^{T}{\mathbf{y}}+{\mathbf{y}}^{T}B\dot{\mathbf{y}}\right)dt\\ \text{where}&\quad A={\mathcal{L}}_{{\mathbf{x}}{\mathbf{x}}},\ B={\mathcal{L}}_{{\mathbf{x}}\dot{\mathbf{x}}},\ C={\mathcal{L}}_{\dot{\mathbf{x}}\dot{\mathbf{x}}}\ ,\end{split}$
and where $\mathbf{y}$} is a differentiable function from $[0,1]$ to ${\mathbb{R}}^{n}$} satisfying homogeneous boundary conditions, ${\mathbf{y}}(0)={\mathbf{y}}(1)=\mathbf{0}$. Neglecting the factor $\frac{1}{2}$ which is irrelevant, we rewrite (15) in the compact form
(16) $\delta^{2}{\mathcal{L}}=\int_{0}^{1}\begin{bmatrix}{\mathbf{y}}^{T}&\dot{\mathbf{y}}^{T}\end{bmatrix}\begin{bmatrix}A&B\\ B^{T}&C\end{bmatrix}\begin{bmatrix}{\mathbf{y}}\\ \dot{\mathbf{y}}\end{bmatrix}dt\ ,\,\ {\mathbf{y}}(0)={\mathbf{y}}(1)=\mathbf{0}\ ,$
and we stress that $A=A^{T}$ and $C=C^{T}$, but in general $B\neq B^{T}$. According to classical results, we need the 2nd variation to be a positive quantity in order for the extremal path ${\mathbf{x}}(t)$ to be a minimizer of the functional $\int_{0}^{1}{\mathcal{L}}({\mathbf{x}},\dot{\mathbf{x}})dt$. The following rewriting is key to obtain conditions guaranteeing positivity of the 2nd variation.
@@
(17) $\begin{split}\delta^{2}{\mathcal{L}}&=\int_{0}^{1}\begin{bmatrix}{\mathbf{y}}^{T}&\dot{\mathbf{y}}^{T}\end{bmatrix}\begin{bmatrix}Q&0\\ 0&P\end{bmatrix}\begin{bmatrix}{\mathbf{y}}\\ \dot{\mathbf{y}}\end{bmatrix}dt\\ \text{where}&\,\ P=P^{T}=C\ ,\,\ Q=Q^{T}=A-\frac{\dot{B}+\dot{B}^{T}}{2}\ .\end{split}$
@@
(18) $\begin{split}&\text{(a)}\quad\dot{\mathbf{y}}^{T}B^{T}{\mathbf{y}}=\frac{d}{dt}({\mathbf{y}}^{T}B^{T}{\mathbf{y}})y^{T}\dot{B}^{T}{\mathbf{y}}-{\mathbf{y}}^{T}B^{T}\dot{\mathbf{y}}\\ &\text{(b)}\quad{\mathbf{y}}^{T}B\dot{\mathbf{y}}=\frac{d}{dt}({\mathbf{y}}^{T}B{\mathbf{y}})-{\mathbf{y}}^{T}\dot{B}{\mathbf{y}}-\dot{\mathbf{y}}^{T}B{\mathbf{y}}\ .\end{split}$
Now, using (LABEL:derivatives)-(a) in (15) and recalling that ${\mathbf{y}}(0)={\mathbf{y}}(1)=0$, we obtain (again, neglecting the factor $\frac{1}{2}$)
$\delta^{2}{\mathcal{L}}=\int_{0}^{1}\left(\dot{\mathbf{y}}^{T}C\dot{\mathbf{y}}+{\mathbf{y}}^{T}(A-\dot{B}^{T}){\mathbf{y}}+{\mathbf{y}}^{T}(B-B^{T})\dot{\mathbf{y}}\right)dt$
@@
(19) $\delta^{2}{\mathcal{L}}=\int_{0}^{1}\begin{bmatrix}{\mathbf{y}}^{T}&\dot{\mathbf{y}}^{T}\end{bmatrix}\begin{bmatrix}A-\dot{B}^{T}&(B-B^{T})/2\\ (B^{T}-B)/2&C\end{bmatrix}\begin{bmatrix}{\mathbf{y}}\\ \dot{\mathbf{y}}\end{bmatrix}dt\ .$
@@
$\delta^{2}{\mathcal{L}}=\int_{0}^{1}\left(\dot{\mathbf{y}}^{T}C\dot{\mathbf{y}}+{\mathbf{y}}^{T}(A-\dot{B}){\mathbf{y}}+\dot{\mathbf{y}}^{T}(B^{T}-B){\mathbf{y}}\right)dt$
@@
(20) $\delta^{2}{\mathcal{L}}=\int_{0}^{1}\begin{bmatrix}{\mathbf{y}}^{T}&\dot{\mathbf{y}}^{T}\end{bmatrix}\begin{bmatrix}A-\dot{B}&(B^{T}-B)/2\\ (B-B^{T})/2&C\end{bmatrix}\begin{bmatrix}{\mathbf{y}}\\ \dot{\mathbf{y}}\end{bmatrix}dt\ .$
Now, adding together (19) and (20), and dividing by $2$, we finally get the expression
(21) $\begin{split}\delta^{2}{\mathcal{L}}&=\int_{0}^{1}\begin{bmatrix}y^{T}&\dot{y}^{T}\end{bmatrix}\begin{bmatrix}A-\frac{\dot{B}+\dot{B}^{T}}{2}&0\\ 0&C\end{bmatrix}\begin{bmatrix}y\\ \dot{y}\end{bmatrix}dt\end{split}$
@@
The expression (17) resembles what people obtain in classic textbooks on calculus of variations ([10, 13, 11]), with the key difference that in these textbooks people restrict to the case $B=B^{T}$, whereas the expression (17) is what we obtain in case $B\neq B^{T}$, and there is the symmetric part of $\dot{B}$ showing up. Recall that $\dot{B}=\frac{d}{dt}{\mathcal{L}}_{{\mathbf{x}}\dot{\mathbf{x}}}$.
With Lemma 3.7, we can now proceed to give verifiable sufficient conditions for having a minimizer. We need $\delta^{2}{\mathcal{L}}$ to be positive when considering all smooth111piecewise smooth would actually suffice functions $\mathbf{y}$} such that ${\mathbf{y}}(0)={\mathbf{y}}(1)=0$. Given the form (17), the following result is well known, see [10], and is given without proof.
@@
Moreover, the absence of conjugate points is equivalent to requiring that the following Riccati equation has a well defined bounded solution over the entire interval $[0,1]$:
(22) $\dot{W}=WP^{-1}W-Q\ ,\quad W(0)=0\ .$
Let us verify these conditions in the cases of interest to us.
@@
As far as the absence of conjugate points, we note that the solution $W$ of (22) (which is symmetric) can be written as follows.
Lemma 3.10.
Let $Y=\begin{bmatrix}U\\ V\end{bmatrix}\in{\mathbb{R}}^{2n,n}$, where $U,V\in{{\mathbb{R}}^{n\times n}}$ are given by the solution of the following linear system:
(23) $\dot{Y}=MY=\begin{bmatrix}0&-P^{-1}\\ -Q&0\end{bmatrix}\begin{bmatrix}U\\ V\end{bmatrix}\ ,\,\ \begin{bmatrix}U\\ V\end{bmatrix}(0)=\begin{bmatrix}I\\ 0\end{bmatrix}\$,
and all blocks are $(n,n)$. Then, for as long as the indicated inverse exists, we have that the solution of (22) is given by
@@
Proof.
The result is well known, and it can be easily verified directly. In fact, that $W(0)=0$ is obvious. Furthermore,
$\frac{d}{dt}V(t)U^{-1}(t)=\dot{V}U^{-1}-VU^{-1}\dot{U}U^{-1}=$
$-QUU^{-1}+VU^{-1}P^{-1}VU^{-1}=-Q+WP^{-1}W\ .$
@@
The matrix in (3.10) is a special case of a Hamiltonian matrix. Recall that a matrix $M=\begin{bmatrix}M_{11}&M_{12}\\ M_{21}&M_{22}\end{bmatrix}$ is Hamiltonian if $M_{12}$ and $M_{21}$ are symmetric and $M_{22}=-M_{11}^{T}$.
Summary 3.12.
For the two costs given by (4), respectively (5), we can solve the Euler-Lagrange equations (12), respectively (14). Based on Lemma 3.5 and Theorem 3.9, in the case of (5), we have reliable criteria to establish both that we have actually satisfied the Euler-Lagrange equation and that the computed solution is a minimizer; we will exploit these facts algorithmically, see Section 4 and 5. However, in the case of cost given by (4), we do not yet have criteria guaranteeing the same.
@@ 3.3. Equivalence between length and energy minimization
Powerful results from Riemannian geometry allow bypassing the difficulties pointed out in Summary 3.12 relative to the cost (4). In fact, the following fundamental result holds; e.g., see [9, Proposition 2.98] or [11, Corollary 2.1.1].
Fact 3.13.
@@
The practical impact of Fact 3.13 is that one does not need to solve the length minimization problem, but solving (14), and verifying that the sufficiency conditions hold, automatically gives the curve of shortest length. In this respect, both approaches give the same geodesic curve. At the same time, it has to be emphasized that the parameterizations of the curve are different; that is, the function $x(t)$, $0\leq t\leq 1$, that solves the Euler-Lagrange (14) does not generally also solve (12).

### v2

@@ 2.2. Sufficient conditions: Conjugate points
As it is well understood, satisfying the Euler-Lagrange equation (9) gives an extremal path, but it is only a necessary condition for a minimizer. The present concern is to obtain sufficient, and verifiable, conditions for having obtained a minimizer in (7). The original work of Gelfand and Fomin, [Gelfand], lays the groundwork for providing sufficient conditions by looking at positivity of the second variation. But unfortunately the derivation of sufficient conditions in [Gelfand] requires the matrix ${\mathcal{L}}_{{\mathbf{x}}\dot{\mathbf{x}}}$ to be symmetric, which is not satisfied in our case; the same assumption is also made in more recent books on the topic (e.g., see [Jost, p.19] and [Liberzon, p.54]). Sufficient conditions without requiring symmetry of ${\mathcal{L}}_{{\mathbf{x}}\dot{\mathbf{x}}}$ are given by Zelikin in [Zelikin], as we will recall below. To formulate these conditions, let us look at the 2nd variation of the functional $\mathcal{L}$} along a trajectory ${\mathbf{x}}(t)$ that solves the Euler-Lagrange equation, in the cases of interest to us. We further elucidates how the sufficient conditions can be constructively verified.
@@
(15) $\begin{split}\delta^{2}{\mathcal{L}}=\frac{1}{2}&\int_{0}^{1}\left({\mathbf{y}}^{T}{\mathcal{L}}_{{\mathbf{x}}{\mathbf{x}}}{\mathbf{y}}+\dot{\mathbf{y}}^{T}{\mathcal{L}}_{\dot{\mathbf{x}}\dot{\mathbf{x}}}\dot{\mathbf{y}}+\dot{\mathbf{y}}^{T}{\mathcal{L}}_{\dot{\mathbf{x}}{\mathbf{x}}}{\mathbf{y}}+{\mathbf{y}}^{T}{\mathcal{L}}_{{\mathbf{x}}\dot{\mathbf{x}}}\dot{\mathbf{y}}\right)dt\\ :=\frac{1}{2}&\int_{0}^{1}\left({\mathbf{y}}^{T}A{\mathbf{y}}+\dot{\mathbf{y}}^{T}C\dot{\mathbf{y}}+\dot{\mathbf{y}}^{T}B^{T}{\mathbf{y}}+{\mathbf{y}}^{T}B\dot{\mathbf{y}}\right)dt\ ,\\ \text{where}&\quad A={\mathcal{L}}_{{\mathbf{x}}{\mathbf{x}}},\ B={\mathcal{L}}_{{\mathbf{x}}\dot{\mathbf{x}}},\ C={\mathcal{L}}_{\dot{\mathbf{x}}\dot{\mathbf{x}}}\ ,\end{split}$
and where $\mathbf{y}$} is a differentiable function from $[0,1]$ to ${\mathbb{R}}^{n}$} satisfying homogeneous boundary conditions, ${\mathbf{y}}(0)={\mathbf{y}}(1)=\mathbf{0}$. We rewrite (15) in the compact form
(16) $\delta^{2}{\mathcal{L}}=\frac{1}{2}\int_{0}^{1}\begin{bmatrix}{\mathbf{y}}^{T}&\dot{\mathbf{y}}^{T}\end{bmatrix}\begin{bmatrix}A&B\\ B^{T}&C\end{bmatrix}\begin{bmatrix}{\mathbf{y}}\\ \dot{\mathbf{y}}\end{bmatrix}dt\ ,\,\ {\mathbf{y}}(0)={\mathbf{y}}(1)=\mathbf{0}\ ,$
and we stress that $A=A^{T}$ and $C=C^{T}$, but in general $B\neq B^{T}$. According to classical results, we need the 2nd variation to be a positive quantity in order for the extremal path ${\mathbf{x}}(t)$ to be a minimizer of the functional $\int_{0}^{1}{\mathcal{L}}({\mathbf{x}},\dot{\mathbf{x}})dt$. The following rewriting of (16) is useful to highlight how, and where, the lack of symmetry of ${\mathcal{L}}_{{\mathbf{x}}\dot{\mathbf{x}}}$ shows up.
@@
(17) $\begin{split}\delta^{2}{\mathcal{L}}&=\frac{1}{2}\int_{0}^{1}\begin{bmatrix}{\mathbf{y}}^{T}&\dot{\mathbf{y}}^{T}\end{bmatrix}\begin{bmatrix}Q&E\\ E^{T}&P\end{bmatrix}\begin{bmatrix}{\mathbf{y}}\\ \dot{\mathbf{y}}\end{bmatrix}dt\ ,\\ \text{where}&\,\ P=P^{T}=C\ ,\,\ Q=Q^{T}=A-\frac{\dot{B}+\dot{B}^{T}}{2}\ ,\,\ E=\frac{B-B^{T}}{2}\ .\end{split}$
@@
(18) $\begin{split}&\text{(a)}\quad\dot{\mathbf{y}}^{T}B^{T}{\mathbf{y}}=\frac{d}{dt}({\mathbf{y}}^{T}B^{T}{\mathbf{y}})-{\mathbf{y}}^{T}\dot{B}^{T}{\mathbf{y}}-{\mathbf{y}}^{T}B^{T}\dot{\mathbf{y}}\ ,\\ &\text{(b)}\quad{\mathbf{y}}^{T}B\dot{\mathbf{y}}=\frac{d}{dt}({\mathbf{y}}^{T}B{\mathbf{y}})-{\mathbf{y}}^{T}\dot{B}{\mathbf{y}}-\dot{\mathbf{y}}^{T}B{\mathbf{y}}\ .\end{split}$
Now, using (LABEL:derivatives)-(a) in (15) and recalling that ${\mathbf{y}}(0)={\mathbf{y}}(1)=0$, we obtain
$\delta^{2}{\mathcal{L}}=\frac{1}{2}\int_{0}^{1}\left(\dot{\mathbf{y}}^{T}C\dot{\mathbf{y}}+{\mathbf{y}}^{T}(A-\dot{B}^{T}){\mathbf{y}}+{\mathbf{y}}^{T}(B-B^{T})\dot{\mathbf{y}}\right)dt\ ,$
@@
(19) $\delta^{2}{\mathcal{L}}=\frac{1}{2}\int_{0}^{1}\begin{bmatrix}{\mathbf{y}}^{T}&\dot{\mathbf{y}}^{T}\end{bmatrix}\begin{bmatrix}A-\dot{B}^{T}&(B-B^{T})/2\\ (B^{T}-B)/2&C\end{bmatrix}\begin{bmatrix}{\mathbf{y}}\\ \dot{\mathbf{y}}\end{bmatrix}dt\ .$
@@
$\delta^{2}{\mathcal{L}}=\frac{1}{2}\int_{0}^{1}\left(\dot{\mathbf{y}}^{T}C\dot{\mathbf{y}}+{\mathbf{y}}^{T}(A-\dot{B}){\mathbf{y}}+\dot{\mathbf{y}}^{T}(B^{T}-B){\mathbf{y}}\right)dt\ ,$
@@
(20) $\delta^{2}{\mathcal{L}}=\frac{1}{2}\int_{0}^{1}\begin{bmatrix}{\mathbf{y}}^{T}&\dot{\mathbf{y}}^{T}\end{bmatrix}\begin{bmatrix}A-\dot{B}&(B^{T}-B)/2\\ (B-B^{T})/2&C\end{bmatrix}\begin{bmatrix}{\mathbf{y}}\\ \dot{\mathbf{y}}\end{bmatrix}dt\ .$
Finally, adding together (19) and (20), and dividing by $2$, we get the expression
(21) $\begin{split}\delta^{2}{\mathcal{L}}&=\frac{1}{2}\int_{0}^{1}\begin{bmatrix}{\mathbf{y}}^{T}&\dot{\mathbf{y}}^{T}\end{bmatrix}\begin{bmatrix}A-\frac{\dot{B}+\dot{B}^{T}}{2}&\frac{B-B^{T}}{2}\\ \frac{B^{T}-B}{2}&C\end{bmatrix}\begin{bmatrix}{\mathbf{y}}\\ \dot{\mathbf{y}}\end{bmatrix}dt\ ,\end{split}$
@@
The expression (17) is what we obtain in case $B\neq B^{T}$, and there is the symmetric part of $\dot{B}$ showing up in $Q$, as well as the antisymmetric part of $B$ in $E$. Recall that $\dot{B}=\frac{d}{dt}{\mathcal{L}}_{{\mathbf{x}}\dot{\mathbf{x}}}$. Of course, in case $B=B^{T}$, then $Q=A-\dot{B}$ and $E=0$.
With Lemma 2.7, we can now proceed to give verifiable sufficient conditions for having a minimizer. We need $\delta^{2}{\mathcal{L}}$ to be positive when considering all smooth111piecewise smooth would actually suffice functions $\mathbf{y}$} such that ${\mathbf{y}}(0)={\mathbf{y}}(1)=0$. The following fundamental result gives the sufficient conditions, see [Gelfand], and is given without proof.
@@
We can verify directly condition (i) of Theorem 2.9, in the cases of interest to us.
@@
As far as the absence of conjugate points, when $P\succ 0$, the absence of conjugate points can be characterized by passing to the Hamiltonian formulation of (17) and from there to the Riccati equation in (23). The underlying process is well understood, and it is the one laid out by Gelfand-Fomin in [Gelfand] and Zelikin in [Zelikin]. We derive it in our case.
Theorem 2.10.
Let $P$ be positive definite. Then, the absence of conjugate points is equivalent to requiring that the $(n,n)$ matrix $U$ in (22) is invertible on $[0,1]$. Here, let $Y=\begin{bmatrix}U\\ V\end{bmatrix}\in{\mathbb{R}}^{2n,n}$, where $U,V\in{{\mathbb{R}}^{n\times n}}$ are given by the solution of the following linear system
(22) $\dot{Y}=MY=\begin{bmatrix}-P^{-1}E^{T}&P^{-1}\\ Q-EP^{-1}E^{T}&EP^{-1}\end{bmatrix}\begin{bmatrix}U\\ V\end{bmatrix}\ ,\,\ \begin{bmatrix}U\\ V\end{bmatrix}(0)=\begin{bmatrix}I\\ 0\end{bmatrix}\$,
and all blocks are $(n,n)$.
Equivalently, the absence of conjugate points is equivalent to requiring that the Riccati equation in (23) has a well defined bounded solution over the entire interval $[0,1]$, where
(23) $\dot{W}=Q-EP^{-1}E^{T}+WP^{-1}E^{T}+EP^{-1}W-WP^{-1}W\ ,\quad W(0)=0\ .$
Proof.
The first step is to take the Euler-Lagrange equation of the functional in the 2nd variation in (17). We have
$\delta^{2}{\mathcal{L}}=\frac{1}{2}\int_{0}^{1}\left(\dot{\mathbf{y}}^{T}P\dot{\mathbf{y}}+\dot{\mathbf{y}}^{T}E^{T}{\mathbf{y}}+{\mathbf{y}}^{T}E\dot{\mathbf{y}}+{\mathbf{y}}^{T}Q{\mathbf{y}}\right)dt\ ,$
and so the Euler-Lagrange equation reads
$(Q{\mathbf{y}}+E\dot{\mathbf{y}})-\frac{d}{dt}(P\dot{\mathbf{y}}+E^{T}{\mathbf{y}})\ =\ 0\ .$
Then, we change variables using the (generalized) momenta ${\mathbf{p}}=P\dot{\mathbf{y}}+E^{T}{\mathbf{y}}$, so that $\dot{\mathbf{y}}=P^{-1}{\mathbf{p}}-P^{-1}E^{T}{\mathbf{y}}$ and from these we obtain the sought linear system (see (22)):
$\begin{cases}\frac{d}{dt}{\mathbf{y}}&=-P^{-1}E^{T}{\mathbf{y}}+P^{-1}{\mathbf{p}}\\ \frac{d}{dt}{\mathbf{p}}&=(Q-EP^{-1}E^{T}){\mathbf{y}}+EP^{-1}{\mathbf{p}}\ .\end{cases}$
With this rewriting, the fact that the absence of conjugate points is equivalent to invertibility of $U$ is a classic result (see [Gelfand, Def.1, Ch.5]).
The equivalence of this fact to the existence of solution to (23) derives from the realization that, for as long as $U$ is invertible, one has that the solution of (23) is given by
@@
This latter result is is easily verified directly. In fact, that $V(0)U^{-1}(0)=0$ is obvious. Furthermore,
$\frac{d}{dt}\bigl(V(t)U^{-1}(t)\bigr)=\dot{V}U^{-1}-VU^{-1}\dot{U}U^{-1}=$
$Q-EP^{-1}E^{T}+(VU^{-1})P^{-1}E^{T}+EP^{-1}(VU^{-1})-(VU^{-1})P^{-1}(VU^{-1})\ ,$
so that $W$ and $VU^{-1}$ satisfy the same differential equation. ∎
@@
The matrix $M$ in (22) is a special case of a Hamiltonian matrix. Recall that a matrix $M=\begin{bmatrix}M_{11}&M_{12}\\ M_{21}&M_{22}\end{bmatrix}$ is Hamiltonian if $M_{12}$ and $M_{21}$ are symmetric and $M_{22}=-M_{11}^{T}$.
Summary 2.12.
For the two costs given by (5), respectively (6), we can solve the Euler-Lagrange equations (12), respectively (14). Based on Lemma 2.5 and Theorem 2.9, in the case of (6), we have reliable criteria to establish both that we have actually satisfied the Euler-Lagrange equation and that the computed solution is a minimizer; we will exploit these facts algorithmically, see Section 3 and 4. However, in the case of cost given by (5), we do not yet have criteria guaranteeing the same.
@@ 2.3. Equivalence between length and energy minimization
Powerful results from Riemannian geometry allow bypassing the difficulties pointed out in Summary 2.12 relative to the cost (5). In fact, the following fundamental result holds; e.g., see [RiemannGeo, Proposition 2.98] or [Jost, Corollary 2.1.1].
Fact 2.13.
@@
The practical impact of Fact 2.13 is that one does not need to solve the length minimization problem, but solving (14), and verifying that the sufficiency conditions hold, automatically gives the curve of shortest length. In this respect, both approaches give the same geodesic curve. At the same time, it has to be emphasized that the parameterizations of the curve are different; that is, the function $x(t)$, $0\leq t\leq 1$, that solves the Euler-Lagrange (14) does not generally also solve (12).